OpenStudy (anonymous):

with the restriction x does not = 2 what is the simplified form of (2x-7/x-2) - (x-5/x-2)

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{b}=\frac{a-c}{b}\] make sure to use parentheses, so the first step is to write \[\frac{2x-7-(x-5)}{x-2}\] then you can clean up the numerator by removing parentheses and combining like terms

OpenStudy (anonymous):

i did that and it gave me x-2/x-2 which is one but that is incorrect

OpenStudy (phi):

for your problem you get =1 (x≠2)

OpenStudy (anonymous):

i got 1 for the simplified answer, but its incorrect, and I don't know why. 2x−7−(x−5) = x - 2 is what i got and x-2/x-2=1

OpenStudy (phi):

what makes you think 1 is wrong ? (strictly speaking the answer is 1 for x≠2) are you sure you typed in the correct question ?

OpenStudy (anonymous):

yes, im sure. I submitted one already, and it was incorrect...

OpenStudy (anonymous):

hmm i am going to stick with that answer

OpenStudy (phi):

(2x-7/x-2) - (x-5/x-2) I assume this is a sloppy version of (2x-7)/(x-2) - (x-5)/(x-2) = 1 for x≠2 however, if we interpret using order of operations 2x - 7/x -2 - (x -5/x -2) which is different. If you are sure the question is (2x-7)/(x-2) - (x-5)/(x-2) then ask your teacher, because the answer is 1, x≠2 see http://www.wolframalpha.com/input/?i=%282x-7%29%2F%28x-2%29+-+%28x-5%29%2F%28x-2%29

OpenStudy (anonymous):

this \[\frac{2x-7}{x-2}-\frac{x-5}{x-2}\] is the correct question right?

OpenStudy (anonymous):

it could be an error in the program, ill try it again though...and yes those are the right questions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hmm...idk, ill ask my teacher. thanks for the help though!

OpenStudy (anonymous):

she says that's correct! thank you both so much!